# Eigenfunction asymptotics in the complex domain for a compact Lie group

### v1

@@ 5.2.2 The Weyl integration formula
[…]
@@ 5.2.3 The proof
$\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,y)$ $\sim$ $\frac{d_{k\,\boldsymbol{\lambda}}}{|W|}\,\int_{\hat{T}}\,\mathrm{d}^{H}V_{\hat{T}}(t)\,\int_{\hat{G}/\hat{T}}\,\mathrm{d}^{H}V_{\hat{G}/\hat{T}}(g\,\hat{T})$
@@
$\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,y)$ $\sim$ $\frac{d_{k\,\boldsymbol{\lambda}}}{|W|}\,\sum_{w\in W}\int_{\hat{T}}\,\mathrm{d}^{H}V_{\tilde{T}}(t)\,\int_{\hat{G}/\hat{T}}\,\mathrm{d}^{H}V_{\hat{G}/\hat{T}}(g\,\hat{T})$
@@
For any $w\in W$, there is $g_{w}\in N(\hat{T})$ (the normalizer of $\hat{T}$ in $\hat{G}$) such that $w(t):=g_{w}\,t\,g_{w}^{-1}$. Furthermore, $\Delta(g_{w}\,t\,g_{w}^{-1})=(-1)^{w}\,\Delta(t)$ by Remark 5.3. We may then rewrite (5.2.3) as follows:
$\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,y)$ $\sim$ $\frac{d_{k\,\boldsymbol{\lambda}}}{|W|}\,\sum_{w\in W}\int_{\hat{T}}\,\mathrm{d}^{H}V_{\hat{T}}(t)\,\int_{\hat{G}/\hat{T}}\,\mathrm{d}^{H}V_{\hat{G}/\hat{T}}(g\,\hat{T})$
@@
$\Psi_{x,y}(u,g\,\hat{T},\boldsymbol{\vartheta})$ ::=:= $u\,\psi^{\tau}\left(\hat{\mu}^{\tau}_{g\,e^{-\imath\,\boldsymbol{\vartheta}}\,g^{-1}}(x),y\right)-\langle\boldsymbol{\lambda},\boldsymbol{\vartheta}\rangle,$ (67)
$\tilde{\mathcal{A}}_{x,y,k}(u,g\,\hat{T},\boldsymbol{\vartheta})$ ::=:= $e^{-\imath\,\langle\boldsymbol{\delta},\boldsymbol{\vartheta}\rangle}\,\rho_{1}\left(g\,e^{\imath\,\boldsymbol{\vartheta}}\,g^{-1},x,y\right)\,s\left(\hat{\mu}^{\tau}_{g\,e^{-\imath\,\boldsymbol{\vartheta}}\,g^{-1}}(x),y,k\,u\right)\,\Delta(e^{\imath\,\boldsymbol{\vartheta}}).$
@@ 5.3.2 The proof
Furthermore, fix $r>0$ sufficiently small (how small may depend on $K$) and let $\rho_{2}\in\mathcal{C}^{\infty}(G)$ be supported in the ball $G_{r}\subseteq G$ centered at $e$ of radius $2\,r$ (say, in the Riemannian metric $\kappa$) and identically equal to $1$ on the ball of radius $r$. Since $\tilde{\mu}^{\tau}$ is free on $K$, the asymptotics of $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,x)$ are unchanged, if the integrand in (58) is multiplied by $\rho_{2}(g)$. Thus
@@
$\boldsymbol{\delta}_{\boldsymbol{\beta}}$ ::=:= $\mathrm{Coad}_{g}(\boldsymbol{\delta})\in\mathcal{O}_{\boldsymbol{\delta}}$ (81)
$L_{k}(\boldsymbol{\beta})$ ::=:= $\mathrm{Coad}_{g}\left(\boldsymbol{\lambda}+\frac{1}{k}\,\boldsymbol{\delta}\right)=\boldsymbol{\beta}+\frac{1}{k}\,\boldsymbol{\delta}_{\boldsymbol{\beta}}\in\mathcal{O}_{\boldsymbol{\lambda}+\frac{1}{k}\,\boldsymbol{\delta}}.$
@@
$\Gamma_{x}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ ::=:= $u\,\psi^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}}}(x),x\right)-\boldsymbol{\beta}(\boldsymbol{\xi})$ (86)
~$\tilde{\mathcal{B}}_{x,k}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ ::=:= $e^{-\boldsymbol{\delta}_{\boldsymbol{\beta}}(\boldsymbol{\xi})}\tilde{\rho}(\boldsymbol{\xi})\,s^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}}}(x),x,k\,u\right)\,\mathcal{V}_{k}(\boldsymbol{\beta})\,\mathcal{P}(\boldsymbol{\xi}),$
@@
The right hand side of (5.3.2) is an oscillatory integral with phase $\Gamma_{x}$ and amplitude ~$\tilde{\mathcal{B}}_{x,k}$. Our next goal is to prove that we may reduce to a compact domain of integration without altering the asymptotics.
@@
$\Upsilon_{x}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ ::=:= $u\,\kappa_{e}\big{(}\mathrm{Ad}_{g}(\boldsymbol{\eta}),\boldsymbol{\xi}\big{)}-\boldsymbol{\beta}(\boldsymbol{\xi})=\left(u\,\tilde{\Phi}^{\tau}(x)-\boldsymbol{\beta}\right)(\boldsymbol{\xi}),$ (96)
$\mathcal{H}_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ ::=:= $e^{-\frac{u}{4\,\tau^{2}}\,\kappa_{e}\big{(}\mathrm{Ad}_{g}(\boldsymbol{\eta}),\boldsymbol{\xi}\big{)}^{2}-\frac{u}{2}\,\|\boldsymbol{\xi}^{\sharp}_{\tilde{G}}(x)\|^{2}+k\,u\,R_{3}\left(\frac{\boldsymbol{\xi}}{\sqrt{k}}\right)}$
@@
To see why Theorem 1.6 follows from Theorem 5.9, recall that in view of the results in [CR1] and (3) one has an a priori uniform bound $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(y,y)\leq C^{\tau}_{\boldsymbol{\lambda}}\,k^{N}$ ($y\in X^{\tau}$) for some $N$ ($N=2d-2$ will do). On the other hand,
@@
Therefore, since $,\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(y,y)$ has at most polynomial growth and $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,x)=O\left(k^{-\infty}\right)$, we also have $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,y)=O\left(k^{-\infty}\right)$ for $k\rightarrow+\infty$. The argument for $P^{\tau}_{k\,\boldsymbol{\lambda}}$ is similar. ∎
@@ 5.4.1 Preliminaries
$\Lambda_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ ::=:= $u\,\psi^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}/\sqrt{k}}}(x_{1k}),x_{2k}\right)-\frac{1}{\sqrt{k}}\,\boldsymbol{\beta}(\boldsymbol{\xi})$ (106)
$\mathcal{D}_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ ::=:= $e^{-\boldsymbol{\delta}_{\boldsymbol{\beta}}(\boldsymbol{\xi})/\sqrt{k}}\,s^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}/\sqrt{k}}}(x_{1k}),x_{2k},k\,u\right)$
@@
$\Psi(u,\boldsymbol{\xi},\boldsymbol{\beta})=\Psi_{\theta_{1},\theta_{2}}(u,\boldsymbol{\xi},\boldsymbol{\beta})$ ::=:= $u\,\left(\theta_{1}-\theta_{2}+\tau\,b_{\boldsymbol{\xi}}\right)-\boldsymbol{\beta}(\boldsymbol{\xi}),$ (114)
@@
::=:= $\imath\,\left[\tilde{\kappa}_{x}\Big{(}{\boldsymbol{\xi}_{\mathfrak{t}}}_{\tilde{G}}(x),{\boldsymbol{\rho}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\rho}_{2}}_{\tilde{G}}(x)\Big{)}+\omega_{x}\Big{(}{\boldsymbol{\xi}_{\mathfrak{s}}}_{\tilde{G}}(x),{\boldsymbol{\eta}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\eta}_{2}}_{\tilde{G}}(x)\Big{)}\right]$
@@
$I_{x,k}(\mathbf{r})$ ::=:= $\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{-\infty}^{+\infty}\,\mathrm{d}b\,\int_{\mathbb{R}^{2\,n_{G}}}\,\mathrm{d}\boldsymbol{\rho}\,\int_{\mathbb{R}^{2\,n_{G}}}\,\mathrm{d}\boldsymbol{\gamma}$ (126)
@@
$-\frac{\imath}{\sqrt{k}}\,L\left(e^{\imath\,\sqrt{k}\,\Psi_{\mathbf{r}}\left(u,b,\boldsymbol{\rho},\boldsymbol{\gamma}\right)}\right)=e^{\imath\,\sqrt{k}\,\Psi_{\mathbf{r}}\left(u,b,\boldsymbol{\rho},\boldsymbol{\gamma}\right)}$
@@
$L_{x}(\mathbf{r})$ ::=:= $\imath\,\tilde{\kappa}_{x}\Big{(}\mathbf{r}_{\tilde{G}}(x),{\boldsymbol{\rho}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\rho}_{2}}_{\tilde{G}}(x)\Big{)}-\frac{1}{2}\,\big{\|}\mathbf{r}_{\tilde{G}}(x)\big{\|}^{2}.$

### v2

@@ 5.2.2 The Weyl integration formula
Let us consider the $\kappa$-orthogonal direct sum $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{t}^{\perp}$. For any $t\in T$, the orthogonal automorphism $\mathrm{Ad}_{t^{-1}}:\mathfrak{g}\rightarrow\mathfrak{g}$ leaves both $\mathfrak{t}$ and $\mathfrak{t}^{\perp}$ invariant; hence, it restricts to an orthogonal endomorphism $(\mathrm{Ad}_{\mathfrak{t}^{\perp}})_{t^{-1}}:\mathfrak{t}^{\perp}\rightarrow\mathfrak{t}^{\perp}$, which doesn’t have the eigenvalue $1$ whenever $t$ is regular. Furthermore, $\det\big{(}(\mathrm{Ad}_{\mathfrak{t}^{\perp}})_{t^{-1}}-\mathrm{Id}_{\mathfrak{t}^{\perp}})>0$ for any $t\in T^{\prime}$.
@@ 5.2.3 The proof
[…]
@@
[…]
@@
[…]
@@
$\Psi_{x,y}(u,g\,\hat{T},\boldsymbol{\vartheta})$ $:=$ $u\,\psi^{\tau}\left(\hat{\mu}^{\tau}_{g\,e^{-\imath\,\boldsymbol{\vartheta}}\,g^{-1}}(x),y\right)-\langle\boldsymbol{\lambda},\boldsymbol{\vartheta}\rangle,$ (68)
$\tilde{\mathcal{A}}_{x,y,k}(u,g\,\hat{T},\boldsymbol{\vartheta})$ $:=$ $e^{-\imath\,\langle\boldsymbol{\delta},\boldsymbol{\vartheta}\rangle}\,\rho_{1}\left(g\,e^{\imath\,\boldsymbol{\vartheta}}\,g^{-1},x,y\right)\,s\left(\hat{\mu}^{\tau}_{g\,e^{-\imath\,\boldsymbol{\vartheta}}\,g^{-1}}(x),y,k\,u\right)\,\Delta(e^{\imath\,\boldsymbol{\vartheta}}).$
@@ 5.3.2 The proof
Furthermore, fix $r>0$ sufficiently small (how small may depend on $K$) and let $\rho_{2}\in\mathcal{C}^{\infty}(G)$ be supported in the ball $G_{r}\subseteq G$ centered at $e$ of radius 2r2\,r (say, in the Riemannian metric $\kappa$) and identically equal to $1$ on the ball of radius $r$. Since $\tilde{\mu}^{\tau}$ is free on $K$, the asymptotics of $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,x)$ are unchanged, if the integrand in (59) is multiplied by $\rho_{2}(g)$. Thus
@@
$\boldsymbol{\delta}_{\boldsymbol{\beta}}$ $:=$ $\mathrm{Coad}_{g}(\boldsymbol{\delta})\in\mathcal{O}_{\boldsymbol{\delta}}$ (82)
$L_{k}(\boldsymbol{\beta})$ $:=$ $\mathrm{Coad}_{g}\left(\boldsymbol{\lambda}+\frac{1}{k}\,\boldsymbol{\delta}\right)=\boldsymbol{\beta}+\frac{1}{k}\,\boldsymbol{\delta}_{\boldsymbol{\beta}}\in\mathcal{O}_{\boldsymbol{\lambda}+\frac{1}{k}\,\boldsymbol{\delta}}.$
@@
$\Gamma_{x}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ $:=$ $u\,\psi^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}}}(x),x\right)-\boldsymbol{\beta}(\boldsymbol{\xi})$ (87)
$\tilde{\mathcal{B}}_{x,k}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ $:=$ $e^{-\boldsymbol{\delta}_{\boldsymbol{\beta}}(\boldsymbol{\xi})}\tilde{\rho}(\boldsymbol{\xi})\,s^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}}}(x),x,k\,u\right)\,\mathcal{V}_{k}(\boldsymbol{\beta})\,\mathcal{P}(\boldsymbol{\xi}),$
@@
The right hand side of (5.3.2) is an oscillatory integral with phase $\Gamma_{x}$ and amplitude $\tilde{\mathcal{B}}_{x,k}$. Our next goal is to prove that we may reduce to a compact domain of integration without altering the asymptotics.
@@
$\Upsilon_{x}(u,\boldsymbol{\beta},\boldsymbol{\xi})$ $:=$ $u\,\kappa_{e}\big{(}\mathrm{Ad}_{g}(\boldsymbol{\eta}),\boldsymbol{\xi}\big{)}-\boldsymbol{\beta}(\boldsymbol{\xi})=\left(u\,\tilde{\Phi}^{\tau}(x)-\boldsymbol{\beta}\right)(\boldsymbol{\xi}),$ (97)
$\mathcal{H}_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ $:=$ $e^{-\frac{u}{4\,\tau^{2}}\,\kappa_{e}\big{(}\mathrm{Ad}_{g}(\boldsymbol{\eta}),\boldsymbol{\xi}\big{)}^{2}-\frac{u}{2}\,\|\boldsymbol{\xi}^{\sharp}_{\tilde{G}}(x)\|^{2}+k\,u\,R_{3}\left(\frac{\boldsymbol{\xi}}{\sqrt{k}}\right)}$
@@
To see why Theorem 1.6 follows from Theorem 5.9, recall that in view of the results in [CR1] and (3) one has an a priori uniform bound $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(y,y)\leq C^{\tau}_{\boldsymbol{\lambda}}\,k^{N}$ ($y\in X^{\tau}$) for some NN (N=2d2N=2d-2 will do). On the other hand,
@@
Therefore, since $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(y,y)$ has at most polynomial growth and $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,x)=O\left(k^{-\infty}\right)$, we also have $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}(x,y)=O\left(k^{-\infty}\right)$ for $k\rightarrow+\infty$. The argument for $P^{\tau}_{k\,\boldsymbol{\lambda}}$ is similar. ∎
@@ 5.4.1 Preliminaries
$\Lambda_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ $:=$ $u\,\psi^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}/\sqrt{k}}}(x_{1k}),x_{2k}\right)-\frac{1}{\sqrt{k}}\,\boldsymbol{\beta}(\boldsymbol{\xi})$ (107)
$\mathcal{D}_{x,k}\left(u,\boldsymbol{\beta},\boldsymbol{\xi}\right)$ $:=$ $e^{-\boldsymbol{\delta}_{\boldsymbol{\beta}}(\boldsymbol{\xi})/\sqrt{k}}\,s^{\tau}\left(\mu^{\tau}_{e^{-\boldsymbol{\xi}/\sqrt{k}}}(x_{1k}),x_{2k},k\,u\right)$
@@
$\Psi(u,\boldsymbol{\xi},\boldsymbol{\beta})=\Psi_{\theta_{1},\theta_{2}}(u,\boldsymbol{\xi},\boldsymbol{\beta})$ $:=$ $u\,\left(\theta_{1}-\theta_{2}+\tau\,b_{\boldsymbol{\xi}}\right)-\boldsymbol{\beta}(\boldsymbol{\xi}),$ (115)
@@
$:=$ $\imath\,\left[\tilde{\kappa}_{x}\Big{(}{\boldsymbol{\xi}_{\mathfrak{t}}}_{\tilde{G}}(x),{\boldsymbol{\rho}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\rho}_{2}}_{\tilde{G}}(x)\Big{)}+\omega_{x}\Big{(}{\boldsymbol{\xi}_{\mathfrak{s}}}_{\tilde{G}}(x),{\boldsymbol{\eta}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\eta}_{2}}_{\tilde{G}}(x)\Big{)}\right]$
@@
$I_{x,k}(\mathbf{r})$ $:=$ $\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{-\infty}^{+\infty}\,\mathrm{d}b\,\int_{\mathbb{R}^{2\,n_{G}}}\,\mathrm{d}\boldsymbol{\rho}\,\int_{\mathbb{R}^{2\,n_{G}}}\,\mathrm{d}\boldsymbol{\gamma}$ (127)
@@
$\frac{\imath}{\sqrt{k}}\,L\left(e^{\imath\,\sqrt{k}\,\Psi_{\mathbf{r}}\left(u,b,\boldsymbol{\rho},\boldsymbol{\gamma}\right)}\right)=e^{\imath\,\sqrt{k}\,\Psi_{\mathbf{r}}\left(u,b,\boldsymbol{\rho},\boldsymbol{\gamma}\right)}$
@@
$L_{x}(\mathbf{r})$ $:=$ $\imath\,\tilde{\kappa}_{x}\Big{(}\mathbf{r}_{\tilde{G}}(x),{\boldsymbol{\rho}_{1}}_{\tilde{G}}(x)+{\boldsymbol{\rho}_{2}}_{\tilde{G}}(x)\Big{)}-\frac{1}{2}\,\big{\|}\mathbf{r}_{\tilde{G}}(x)\big{\|}^{2}.$